OpenStudy (kittiwitti1):

SINGLE-VAR CALCULUS whole problem: "Find an equation to the tangent line to the curve at the given point" with given equation - y=log2(x^2-3x+6) at (2,2) Step one - derivative of:\[y=\log{2}(x^{2}-3x+6)\] I thought I'd use the Product Rule but Symbolab tells me to leave the \(\log{2}\) alone. :\

OpenStudy (afro):

why would you lead him on, WHY WOULD YOU LEAD THIS WHOLE THING ON?

563blackghost (563blackghost):

What is the question?

OpenStudy (kittiwitti1):

@Zale101

OpenStudy (3mar):

\[y=\log_2(x^2-3x+6)\]???

OpenStudy (kittiwitti1):

log(2) not log-base 2

OpenStudy (3mar):

\[y=\log2(x^2-3x+6)\]

OpenStudy (kittiwitti1):

Yes, that.

OpenStudy (3mar):

Here is the rule.

OpenStudy (3mar):

Could you apply the rule of differentiation ?

OpenStudy (kittiwitti1):

What do you mean?

OpenStudy (3mar):

Of course you know that \[\log(ax)=\log a+\log (x)\] and due to log a is a constant, it gets out of the derivaitve of the main function

OpenStudy (3mar):

for ex: \[\frac{ d }{ dx }(2x^2)=2\frac{ d }{ dx }(x^2)\]

OpenStudy (kittiwitti1):

I would apply the product rule theoretically but Symbolab told me to leave as is. (My thought process)\[\frac{d}{dx}\left((\log{2})\times(x^{2}-3x+6)\right)\]and apply the Product Rule from there

OpenStudy (kittiwitti1):

Yet what I was told was to leave off the \(\log{2}\) as \(\log_{10}{2}\)

OpenStudy (3mar):

OOOOOOOOOOOOOOOHHHHHHHHHHHHHHH You wrote it wrong,MAN This is not fair. Why you did not write it correctly?

OpenStudy (kittiwitti1):

I did?

OpenStudy (kittiwitti1):

The problem on my paper is written as\[\log{2}(x^{2}-3x+6)\]

OpenStudy (kittiwitti1):

Here is the entire Symbolab result: https://goo.gl/Y3Xlzx

OpenStudy (3mar):

You mean that \[\frac{ d }{ dx }[(\log2)(x^2-3x+6)]=\frac{ d }{ dx }[constant*polynimial]\] \[=constant\frac{ d }{ dx }[polynomial]\]

OpenStudy (3mar):

@kittiwitti1

OpenStudy (loser66):

Question: the tangent passes through (2,2) or tangent AT (2,2) ??

OpenStudy (kittiwitti1):

@3mar eh? @Loser66 "Find an equation to the tangent line to the curve at the given point"

OpenStudy (loser66):

and like @3mar question. I would like to know which one is correct one \(y =log (2x^2-6x+12)\) (1) or \(y= (x^2-3x+6) log2\) (2) Which one??

OpenStudy (3mar):

(2)

OpenStudy (loser66):

so, it is simple enough. :)

OpenStudy (kittiwitti1):

I don't know. That's what the paper says.

OpenStudy (3mar):

I told you!

OpenStudy (kittiwitti1):

?

OpenStudy (loser66):

but!!

OpenStudy (kittiwitti1):

I applied the Product Rule as multiplying two forms together.

OpenStudy (loser66):

(2,2) is not on the curve, how can you find the tangent at (2,2)?

OpenStudy (3mar):

You don't need to apply the Product Rule. Just derive the polynomial

OpenStudy (loser66):

Passes through (2,2) , we have a way to find it out.

OpenStudy (kittiwitti1):

Derive polynomial? @3mar

OpenStudy (3mar):

Yes. It is simple now

OpenStudy (kittiwitti1):

I think you are right @Loser66 the tangent line PASSES through (2,2)

OpenStudy (kittiwitti1):

OH, because it is not \(\log_{b}{x}\) but rather\(\log_{b}{a}\), yes?

OpenStudy (kittiwitti1):

Thus not a function of \(x\)

OpenStudy (3mar):

\[\frac{ dy }{ dx }=(\log2)*\frac{ dy }{ dx }(x^2-3x+6)=\log2(2x-3)\] apply (2,2) \[\frac{ dy }{ dx }|_{x=2}=\log2(2*2-3)=(\log2)*(1)=\log2=0.301\]

OpenStudy (3mar):

Yes you are right

OpenStudy (kittiwitti1):

Why does the constant stay outside? Isn't derivative of a constant 0... D: I'm confused

OpenStudy (3mar):

NO MAN Because it is multiplied by the polynomial \[\frac{ d }{ dx }(constant*x^2)=constant*\frac{ d }{ dx }(x^2)\]

OpenStudy (3mar):

log 2 is constant not a function multiplied by a function.

OpenStudy (kittiwitti1):

But it is within the y-function (equation?) o.o

OpenStudy (3mar):

Y-function may contain constants in the polynomial. ex: \[y=x^2+12x+36\] 16 is a constant

OpenStudy (kittiwitti1):

16?\[y'=\left(x^{2}+12x+36\right)'\rightarrow(2x+12(1)+0)=2x+12\]

OpenStudy (kittiwitti1):

I get the feeling I am applying the wrong logic to this

OpenStudy (3mar):

You got it!

OpenStudy (kittiwitti1):

But that would make log 2 = 0 by derivative constant rule ... ?

OpenStudy (kittiwitti1):

... I DO apply the Product Rule, don't I :\

OpenStudy (3mar):

"But that would make log 2 = 0 by derivative constant rule ... ?" This is in case of dy/dx(log2+x^2-34+6) not dy/dx((log 2)(x^2-34+6)) got it?

OpenStudy (kittiwitti1):

\[\left(\log{2}(x^{2}-3x+6)\right)'=(\log{2})'•(x^{2}-3x+6)+\log{2}\left(x^{2}-3x+6\right)'\]\[=0(x^{2}-3x+6)+\log{2}(2x-3(1)+0)=0+\log2(2x-3)=\log2\times(2x-3)\]

OpenStudy (kittiwitti1):

Does that work? D:

OpenStudy (3mar):

Yes. This is the applying the rule literally. Great job! we can deduce from this the rule: \[\frac{ d }{ dx }(5x^3)=5\frac{ d }{ dx }(x^3)=5(3x^2)\]

OpenStudy (kittiwitti1):

Oh, okay. I was utterly confused for the longest time lol Thank you! @3mar

OpenStudy (3mar):

But now?

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (kittiwitti1):

I'm good now. ☺

OpenStudy (3mar):

Thanks for Allah I am happy to hear that!

OpenStudy (3mar):

Thank you for the medal!